OpenStudy (nuccioreggie):

What will be the perimeter and the area of the rectangle below if it is enlarged using a scale factor of 4.5? A rectangle is shown. The length of the rectangle is labeled as 8 cm, and the width is labeled as 6 cm. Perimeter = 46 cm, area = 131.25 cm2 Perimeter = 126 cm, area = 972 cm2 Perimeter = 46 cm, area = 972 cm2 Perimeter = 126 cm, area = 131.25 cm2

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (nuccioreggie):

i think its B

OpenStudy (anonymous):

its B

OpenStudy (nuccioreggie):

@Michele_Laino. Waht is it

OpenStudy (michele_laino):

the new dimensions are: 8*4.5=... 6*4.5=...

OpenStudy (nuccioreggie):

36 n 3

OpenStudy (nuccioreggie):

Is the correct answr B

OpenStudy (welshfella):

perimeter = 2*27 + 2 * 36 area = 27*36

OpenStudy (welshfella):

because width = 4.5 * 6 = 27 and length = 8 * 4.5 = 36

OpenStudy (nuccioreggie):

So is the answer B or NO????????

OpenStudy (anonymous):

ITS B

OpenStudy (nuccioreggie):

How do u know explain